# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v1

@@ Acknowledgements.
We begin by explaining the procedure for generating the $A$ matrix of size $n$. We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [26] to construct the Hamiltonian for a given molecule in a specified single particle basis. GUGA-CI ensures that the Hamiltonian is constructed in an orthogonal basis; hence, it is Hermitian. Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results. We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature). The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the Hartree-Fock (HF) energy and the correlation energy, with the latter being much lesser than the former). This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself. We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite. In a sense, subtracting MP2 energy could be thought of as a level shift over and beyond subtracting the HF energy.
@@
$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}.$ (A25)
@@
Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b) and (c) show that the $A$ matrix increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
@@
We fix the number of core orbitals to 13. We now describe the method adopted in choosing virtual molecular orbitals (MOs). Instead of taking the approach of ordering the virtual MOs by energy in ascending order and then imposing an energy cut-off to define our active space, we pick those MOs that are dominated by important atomic orbitals (AOs). We select the important AOs, and for the purposes of this work, we pick those AOs in the order of the principal quantum numbers. For example, in the case of RbH, we can decide the important AOs from the set that arise from our choice of basis, namely the Sapporo double zeta basis: $5s$ through $7p$. For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A9). We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 13 occupied MOs and 1 virtual one has a system size, $m$, of $40\times 40$, and subsequently upon padding, an $A$ matrix size, $n$, of $64\times 64$. For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5p_{y}$, leading to a $119\times 119$ matrix, which, after padding, becomes a matrix of size $128\times 128$. We continue to construct $A$ matrices in this fashion until we reach $m=9065$, for which $\mathcal{\kappa}$ is about 348.
@@
For the case of CsH molecule, we fix the number of occupied orbitals to 18 while progressively increasing the number of valence orbitals. The order in which we add virtual MOs as well as each such virtual MO’s details are mentioned in Table A9.
Refer to caption Refer to caption
(a) (b)
Figure A10: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the RbH molecule (sub-figure (a)) and the CsH molecule (sub-figures (b)) in the Sapporo double zeta basis set. The system size is determined by increase in number of virtuals.
Refer to caption Refer to caption
(a) (b)
Figure A11: Plots showing the increase in sparsity, s, with $A$ matrix size, $m$, for the RbH (sub-figure (a)) and the CsH (sub-figure (b)) molecules in the Sapporo double zeta basis set. The system size is determined by increase in number of virtuals.
@@
Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $n\times n$ $A$ matrix versus system size ($n$), for RbH and CsH molecules, where we vary the virtual orbitals to increase $n$. In the table, $n_{o}$ the number of active spin orbitals, and $n_{v}$ the number of unoccupied spin orbitals. The first row of the column, ‘MO character’, gives information on the atomic orbitals that contribute dominantly to the considered virtual molecular orbital for $m=40$ case, whereas the second row gives the same information about the second virtual added to the active space for $m=119$ case, and so on. The last column specifies the corresponding newly added virtual’s energy in units of Hartree.
Molecule $m$ $\mathcal{\kappa}(n_{o},n_{v})$ MO character Virtual orbital energy (in Hartree)
RbH 40 152.23 (13, 1) Rb $5s-5p_{z}$ hybrid -0.02
119 152.91 (13, 2) Rb $5p_{y}$ 0.04
228 153.18 (13, 3) Rb $5p_{x}$ 0.04
407 157.92 (13, 4) Rb $5s-5p_{z}$ hybrid 0.08
889 192.30 (13, 6) Rb $4d_{x^{2}-y^{2}}$, Rb $4d_{xy}$ 0.57, 0.57
1863 195.79 (13, 9) Rb $4d_{xz}$, Rb $4d_{yz}$, Rb $4d_{z^{2}}$ 0.57, 0.57
3866 228.12 (13, 13) Rb $6s$, Rb $6p_{x}$, Rb $6p_{y}$, Rb $6p_{z}$ 0.57, 0.06, 0.42, 0.42
4518 235.94 (13, 14) Rb $7s$ 0.44
5225 312.02 (13, 15) H $2s$ 0.83
5896 336.53 (13, 16) H $2p_{y}$ 0.87
6595 338.98 (13, 17) H $2p_{x}$ 1.69
7441 345.62 (13, 18) H $2p_{z}$ 1.69
8238 346.55 (13, 19) Rb $7p_{y}$ 1.71
9065 347.49 (13, 21) Rb $7p_{x}$, Rb $7p_{z}$ 2.12, 2.14
CsH 67 465.89 (18, 1) Cs $6p_{x}$ 0.42
210 466.59 (18, 2) Cs $6p_{y}$ 0.42
417 467.08 (18, 3) Cs $6p_{z}$ 0.48
744 487.04 (18, 4) Cs $6s$ 1.49
1171 541.58 (18, 5) Cs $8s$ 0.01
1698 546.70 (18, 6) Cs $7p_{y}$ 0.05
2175 551.19(18, 7) Cs $7p_{x}$ 0.05
2802 573.09(18, 8) Cs $7s$ 0.08
3493 593.31(18, 9) Cs $5d_{y^{2}}$ 0.38
4344 640.22 (18, 10) Cs $5d_{xy}$ 0.38
5215 646.10 (18, 11) Cs $5d_{xz}$ 0.39
6150 652.10 (18, 12) Cs $5d_{yz}$ 0.39
7261 714.59 (18, 13) Cs $5d_{z^{2}}$ 0.39
8473 956.37 (18, 14) H $2s$ 0.93
9784 974.45 (18, 15) H $2p_{x}$ 1.70
11058 1003.49 (18, 16) H $2p_{y}$ 1.70
@@
Table A10: Table providing the minimum and maximum eigenvalues of the matrix $A_{mol}$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$, for RbH and CsH systems. The energies are in units of milli-Hartree.
Molecule $m$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
RbH 40 135.31 20599.95
119 135.18 20670.59
228 135.04 20685.98
407 131.94 20837.37
889 109.40 21038.25
1863 107.55 21058.36
3866 92.36 21071.00
4518 90.25 21294.40
5225 72.72 22690.04
5896 72.32 24341.29
6595 71.93 24386.59
7441 71.46 24701.91
8238 71.27 24702.28
9065 71.09 24702.65
CsH 67 128.09 59680.83
210 127.99 59722.33
417 127.90 59737.34
744 122.88 59847.85
1171 110.54 59867.65
1698 109.50 59867.64
2175 108.61 59867.63
2803 104.46 59867.67
3493 100.89 59867.69
4345 93.60 59916.79
5215 92.75 59929.31
6150 91.91 59939.41
7262 83.88 59948.78
8473 64.56 61759.13
9784 63.38 61761.97
11058 63.22 63448.99
@@
Table A11: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2, RbH and CsH molecules resp. in STO-6G, Sapporo-DZP and Sapporo-DKH3-DZP-2012 basis sets. The PFDs are calculated over a range of 5 repetitions.
Li2 RbH CsH
# shots HHL(PFD, Range) Psi-HHL(PFD, Range) HHL(PFD, Range) Psi-HHL(Range) HHL(PFD, Range) Psi-HHL(PFD, Range)
@@
$10^{3}$ (-4.08, 19.80) (-1.46, 8.34) (1.15, 30.41) (2.51, 25.93) (-3.76, 55.34) (0.49, 23.56)
$5\times 10^{3}$ (-2.94, 8.89) (-2.11, 7.57)
$10^{4}$ (-0.58, 6.00) (1.07, 4.47) (3.64, 25.05) (-0.42, 6.18) (-11.83, 9.82) (2.24, 4.22)
$5\times 10^{4}$ (3.29, 16.90) (1.92, 3.14) (-12.38, 2.04) (1.63, 4.77)
$10^{5}$ (5.77, 4.03) (0.60, 1.79) (-11.74, 2.49) (-0.06, 2.29)
$5\times 10^{5}$ (5.29, 0.76) (0.99, 0.71) (-10.49, 2.29) (0.38, 1.77)
$10^{6}$ (5.46, 0.72) (1.07, 0.56) (-10.92, 2.01) (0.56, 1.05)

### v2

@@ Acknowledgements.
We begin by explaining the procedure for generating the $A$ matrix of size $n$. We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [27] to construct the Hamiltonian for a given molecule in a specified single particle basis. GUGA-CI ensures that the Hamiltonian is constructed in an orthogonal basis; hence, it is Hermitian. Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results. We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature). The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much lesser than the former). This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself. We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite. In a sense, subtracting MP2 energy could be thought of as a level shift over and beyond subtracting the HF energy.
@@
$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}.$ (A25)
@@
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that the $A$ matrix of respectively KH, RbH and CsH increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
@@
We fix the number of core orbitals to 6, 8 and 10. We now describe the method adopted in choosing virtual molecular orbitals (MOs). Instead of taking the approach of ordering the virtual MOs by energy in ascending order and then imposing an energy cut-off to define our active space, we pick those MOs that are dominated by important atomic orbitals (AOs). We select the important AOs, and for the purposes of this work, we pick those AOs in the order of the principal quantum numbers. For example, in the case of RbH, we can decide the important AOs from the set that arise from our choice of basis, namely the Sapporo double zeta basis: $5s$ through $7p$. For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A9). We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n$, of $256\times 256$. For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5s$, leading to a $279\times 279$ matrix, which, after padding, becomes a matrix of size $512\times 512$. We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 389.76.
@@
We do the same for KH and CsH molecules, we fix the number of occupied orbitals to 6, 8 and 10 occupied oritals while progressively increasing the number of valence orbitals. The order in which we add virtual MOs as well as each such virtual MO’s details are mentioned in Table A9.
@@
Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$. In the table, $n_{o}$ is the number of active spin orbitals, and $n_{v}$ is the number of unoccupied spin orbitals. The column labelled ‘Reordered virtual MOs’ represents the atomic orbitals reordered according to the ones that contribute dominantly to the virtual molecular orbital. The column labelled ‘MO Energies’ gives the molecular virtual orbital energies in units of milli-Hartree.
Molecule Reordered virtual MOs MO Energies $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$
K $4s$ 1151.70 235 233.65(6, 6) 580 41.03(8, 8) 1484 2195.80(10, 10)
K $4p_{y}$ 3609.70 313 234.44(6, 7) 745 41.57(8, 9) 1761 2206.59(10, 11)
K $4p_{x}$ 3609.70 370 235.12(6, 8) 932 236.57(8, 10) 2046 2217.42(10, 12)
K $4p_{z}$ 3663.70 432 236.80(6, 9) 1110 237.74(8, 11) 2451 2323.69(10, 13)
K $3d_{xy}$ 743.00 541 238.28(6, 10) 1296 238.90(8, 12) 2900 2849.98(10, 14)
K $3dy^{2}$ 743.00 630 244.66(6, 11) 1540 250.34(8, 13) 3393 2904.40(10, 15)
K $3d_{yz}$ 747.40 748 250.45(6, 12) 1808 307.03(8, 14) 3794 2933.74(10, 16)
K $3d_{xz}$ 747.40 889 265.31(6, 13) 2100 312.83(8, 15) 4203 2964.04(10, 17)
K $3dz^{2}$ 755.30 1044 340.59(6, 14) 2358 316.09(8, 16) 4788 3610.73(10, 18)
K $6s$ 3.70 1213 348.68(6, 15) 2624 319.44(8,17) 5265 3629.57(10,19)
K $6p_{y}$ 72.70 1338 371.54(6, 16) 2972 389.11(8,18) 5750 3648.68(10,20)
K $6p_{x}$ 72.70 1510 374.07(6, 17) 3278 391.14(8, 19) 6427 3674.83(10, 21)
K $6p_{z}$ 82.80 1711 380.25(6, 18) 3592 393.20(8, 20)
K $5s$ 147.20 1860 385.51(6, 19) 3996 395.77(8, 21)
K $5p_{z}$ 692.70 2062 386.30(6, 20)
K $5p_{y}$ 702.90 2296 389.76(6, 21)
K $5p_{x}$ 702.90
H $2s$ 934.00
H $2p_{y}$ 1668.70
H $2p_{x}$ 1668.70
KH H $2p_{z}$ 1914.10
Rb $5s-5p_{z}$ 940.00 209 102.73(6, 6) 632 104.88(8, 8) 1401 108.16(10,10)
Rb $5s$ 925.80 279 103.00(6, 7) 744 105.81(8, 9) 1673 111.85(10, 11)
Rb $5p_{x}$ 2283.00 370 103.74(6, 8) 932 106.63(8, 10) 1963 115.27(10, 12)
Rb $5p_{y}$ 2283.00 432 104.65(6, 9) 1110 110.27(8, 11) 2321 125.33(10, 13)
Rb $5p_{z}$ 2305.80 541 105.46(6, 10) 1296 113.64(8, 12) 2713 156.32(10, 14)
Rb $4d_{xy}$ 635.40 630 109.05(6, 11) 1540 123.51(8, 13) 3141 182.23(10, 15)
Rb $4dx^{2}$ 635.40 748 112.38(6, 12) 1808 154.29(8, 14) 3537 206.22(10, 16)
Rb $4d_{xz}$ 641.20 889 122.12(6, 13) 2100 179.90(8, 15) 3951 207.93(10, 17)
Rb $4d_{yz}$ 641.20 1044 152.70(6, 14) 2358 203.86(8, 16) 4460 213.88(10, 18)
Rb $4dz^{2}$ 666.90 1213 178.09(6, 15) 2624 205.55(8, 17) 4932 214.42(10, 19)
Rb $7s$ 3.20 1381 202.02(6, 16) 2972 211.53(8, 18) 5422 214.96(10, 20)
Rb $7p_{x}$ 64.20 1510 203.70(6, 17) 3278 212.06(8, 19) 6013 216.27(10, 21)
Rb $7p_{y}$ 64.20 1711 209.67(6, 18) 3592 212.60(8, 20)
Rb $6s$ 78.40 1860 210.19(6, 19) 3996 213.87(8, 21)
Rb $6s$ 130.80 2062 210.72(6, 20)
Rb $6p_{x}$ 477.70 2296 211.94(6, 21)
Rb $6p_{y}$ 477.70
Rb $6p_{z}$ 487.50
H $2p_{x}$ 1678.20
[…]
Table A10: (Continued.)
Molecule Reordered virtual MOs MO Energies $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$
Cs $6p_{x}$ 364.80 209 77.05(6, 6) 632 81.29(8, 8) 1484 94.54(10, 10)
Cs $6p_{y}$ 364.80 279 77.18(6, 7) 812 92.18(8, 9) 1711 95.68(10, 11)
Cs $6p_{z}$ 382.00 370 80.54(6, 8) 1016 93.40(8, 10) 2007 100.02(10, 12)
Cs $6s$ 1455.30 475 91.32(6, 9) 1044 94.51(8, 11) 2321 104.14(10, 13)
Cs $8s$ 172.00 595 92.54(6, 10) 1339 98.79(8, 12) 2714 114.07(10, 14)
[…]
Cs $7p_{x}$ 25.50 759 97.83(6, 12) 1808 112.65(8, 14) 3537 184.86(10, 16)
Cs $7s$ 75.60 890 101.83(6, 13) 2100 156.01(8, 15) 3951 186.03(10,17)
Cs $7s$ 85.50 1044 111.49(6, 14) 2358 182.96(8, 16) 4461 192.15(10, 18)
Cs $5dy^{2}$ 307.30 1213 154.45(6, 15) 2624 184.12(8, 17)
Cs $5d_{xy}$ 307.30 1338 181.34(6, 16) 2973 190.17(8, 18)
Cs $5d_{xz}$ 309.70 1510 182.49(6, 17)
Cs $5d_{yz}$ 309.70 1711 188.52(6, 18)
Cs $5dz^{2}$ 310.90
H $2s$ 882.50
H $2p_{x}$ 1716.80
H $2p_{y}$ 1716.80
CsH H $2p_{z}$ 1716.80
@@
Table A11: Table providing the minimum and maximum eigenvalues of the matrix $\mathrm{A}_{mol}$ of size $m$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ respectively, for KH, RbH and CsH systems. $n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals respectively. The energies are in units of milli-Hartree.
Molecule $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
202 6 733.73 25918.07 580 8 730.05 29960.23 1484 10 127.91 280908.57
274 7 731.91 29570.15 745 9 728.20 30271.74 1761 11 127.29 280909.73
337 8 730.17 29604.25 932 10 127.95 30272.05 2046 12 126.64 280910.95
432 9 728.35 29727.55 1110 11 127.33 30272.33 2451 13 120.90 280914.12
541 10 128.01 29729.71 1296 12 126.71 30272.59 2900 14 98.56 280914.37
655 11 127.38 29730.03 1540 13 120.92 30272.79 3393 15 96.78 281150.00
748 12 126.76 29730.35 1808 14 98.59 30272.81 3794 16 95.80 281161.10
889 13 120.98 29731.63 2100 15 96.83 30293.88 4203 17 94.84 281171.57
1043 14 98.65 29731.69 2358 16 95.87 30304.25 4788 18 77.86 281200.60
1213 15 96.89 29830.59 2624 17 94.89 30314.00 5265 19 77.45 281200.93
1381 16 95.93 29834.52 2972 18 77.91 30317.70 5750 20 76.99 281201.42
1510 17 94.96 29839.08 3278 19 77.50 30317.87 6427 21 76.56 281469.15
1711 18 77.98 29854.11 3592 20 77.10 30318.05
1909 19 77.57 29854.20 3996 21 76.62 30328.12
2062 20 77.17 29854.30
KH 2295 21 76.69 29894.66
209 6 108.95 11193.43 632 8 108.02 11330.03 1401 10 106.22 11489.85
279 7 108.73 11200.46 744 9 107.07 11330.11 1673 11 102.72 11490.43
370 8 108.02 11206.98 932 10 106.25 11330.19 1963 12 99.68 11491.02
432 9 107.08 11207.05 1110 11 102.75 11330.84 2321 13 91.77 11503.07
541 10 106.26 11207.13 1296 12 99.70 11331.49 2713 14 81.20 12694.32
630 11 102.77 11207.95 1540 13 91.80 11339.56 3141 15 72.02 13125.74
748 12 99.73 11208.41 1808 14 81.23 12533.58 3537 16 71.65 14777.36
889 13 91.84 11216.40 2100 15 72.05 12963.58 3951 17 71.28 14822.69
1044 14 81.27 12410.75 2358 16 71.68 14614.13 4460 18 70.82 15148.61
1213 15 72.10 12840.73 2624 17 71.31 14659.47 4932 19 70.64 15148.86
1381 16 71.72 14490.85 2972 18 70.85 14988.70 5422 20 70.46 15149.12
1510 17 71.35 14536.14 3278 19 70.67 14988.95 6013 21 70.05 15151.35
1711 18 70.90 14865.97 3592 20 70.50 14989.21
1860 19 70.72 14866.22 3996 21 70.09 14991.44
2062 20 70.55 14866.47
RbH 2295 21 70.15 14868.69
209 6 119.60 9216.72 632 8 114.40 9306.7 1484 10 99.94 9447.99
279 7 119.41 9216.73 812 9 101.3 9338.37 1711 11 98.74 9448.39
370 8 114.52 9224.27 1016 10 99.97 9338.71 2007 12 94.46 9448.85
475 9 101.35 9255.84 1044 11 98.81 9339.05 2321 13 90.73 9449.29
595 10 100.02 9256.34 1339 12 94.53 9339.23 2714 14 82.84 9449.94
664 11 98.86 9256.39 1540 13 90.80 9339.41 3141 15 63.54 10033.15
759 12 94.61 9256.47 1808 14 82.91 9340.27 3537 16 63.38 11718.57
890 13 90.89 9256.59 2100 15 63.60 9923.04 3951 17 63.23 11763.91
1044 14 83.03 9257.34 2358 16 63.44 11609.02 4461 18 62.86 12079.72
1213 15 63.71 9840.84 2624 17 63.29 11654.36
1338 16 63.56 11526.12 2973 18 62.92 11966.65
1510 17 63.40 11571.52
CsH 1711 18 63.04 11884.73
@@
Table A12: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The PFDs are calculated over a range of 30 repetitions.
Li2 KH
# shots HHL(PFD, Range) Psi-HHL(PFD, Range) HHL(PFD, Range) Psi-HHL(PFD, Range)
@@
$10^{3}$ (-4.08, 19.80) (-1.46, 8.34) (nan, nan) (-11.98, 106.87)
$5\times 10^{3}$ (-2.94, 8.89) (-2.11, 7.56)
$10^{4}$ (-0.58, 6.00) (1.07, 4.47) (0.34, 68.36) (1.29, 44.70)
$5\times 10^{4}$ (0.73, 32.51) (0.68, 16.78)
$10^{5}$ (0.73, 17.00) (0.72, 9.02)
$5\times 10^{5}$ (2.65, 10.59) (1.11, 5.65)
$10^{6}$ (3.99, 5.09) (0.22, 4.05)
Table A13: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the LCCSD total energy and the PFDs over 30 repetitions with respect to the bond length of KH molecule. All the energies are in units of milli-Hartree.
Bond length (Bohr) $\kappa$ HHL Psi-HHL
Total energy $\Delta E$ PFD Total energy $\Delta E$ PFD
2.82 393.21 -598878.025 -0.124 -3.92 -598877.968 -0.068 -2.48
3.29 374.93 -598940.871 -0.086 -3.84 -598940.844 -0.059 2.31
3.76 367.32 -598967.447 -0.003 -1.96 -598967.447 -0.004 -0.26
4.23 365.40 -598975.800 0.007 1.46 -598975.802 0.004 0.34
4.71 371.38 -598974.517 -0.010 -10.00 -598974.509 -0.007 1.39
5.18 385.16 -598968.421 -0.001 -0.44 -598968.419 $0$0.08
5.65 373.56 -598960.199 -0.001 -0.41 -598960.196 0.000 0.26
Figure A10: Figure showing average PFDs and range of PFDs (over 20 repetitions) for different values of $a$ for the overlap, $|\langle 0|(a|0\rangle+b|1\rangle)|$. We fix the number of shots to 1000 for generating all of the data.